For the points (-1, 1) and (2, 5), find the distance d between the points and the midpoint M of the line segment which connects them.
Found 2 solutions by josgarithmetic, rothauserc:
Answer by josgarithmetic(39617)

Distance Formula, and Midpoint Formula

Answer by rothauserc(4718)

d = square root((2-(-1)^2 +(5-1)^2) = 5
:
M = ((2+(-1)/2), (5+1)/2) = (1/2, 3)
: